The diameter of the Sun is 1.4x 106 kilometers. (1 km = 3280 ft) What is the volume of the Sun in cubic feet?
Answer #1

Volume of sphere is given by:

V = 4*pi*r^3/3

r = radius of sphere

So Using given values:

r = radius of sun = diameter/2

r = 1.4*10^6 km/2 = 0.7*10^6 km

Given that 1 km = 3280 ft

r = 0.7*10^6 km*(3280 ft/1 km) = 2.296*10^9 ft

So Volume of sun will be

V = 4*pi*(2.296*10^9)^3/3

V = 5.07*10^28 ft^3

Please Upvote.

Comment below if you have any query.